The index of refraction for a certain type of glas
Answer #1

The Snell's Law express that:

n_{1}.sin(\theta _{1})=n_{2}.sin(\theta _{2})

Solving for \theta _{2} for both wavelengths (blue and red) using their refraction index n2, we have:

\theta _{2Blue}=sin^{-1}(\frac{sin(35,05)}{1,643})=20,459

\theta _{2Red}=sin^{-1}(\frac{sin(35,05)}{1,605})=20,966

Finally, the difference between angles is:

\left | \theta _{2}-\theta _{1} \right |=\left |20,966- 20,459 \right |=0,507